OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (2, 5) and (3, 0). y − = (x − 2)

OpenStudy (anonymous):

!!

OpenStudy (solomonzelman):

First, find the slope (between the 2 points). Do you know how to do this?

OpenStudy (solomonzelman):

\(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2}}{\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2}} }\) Where, 1. \(\LARGE \color{black}{ \displaystyle {\rm m} }\) is the slope 2. \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\)

OpenStudy (solomonzelman):

use this formula.

OpenStudy (anonymous):

no i dont know how to

OpenStudy (solomonzelman):